Clue #1: The theme of the cache is "Superheroes." Carefully graph the four lines with the
following equations on the coordinate plane below to see the shape of the item.
A. y=2x+13
B. 2x-y=17
C. y=-1/2x+13
D. x+2y=-14
My question is how do I figure out the x and y values of these equations?

Answer by Alan3354(69443)

Clue #1: The theme of the cache is "Superheroes." Carefully graph the four lines with the
following equations on the coordinate plane below to see the shape of the item.
A. y=2x+13
B. 2x-y=17
C. y=-1/2x+13
D. x+2y=-14
My question is how do I figure out the x and y values of these equations?
====================
A. y=2x+13
Pick any 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.
That's all it takes.